Find the 9th roots of unity. Find the 6th roots of -64.
Answer #1

a.) To find the 9th roots of unity:

We have,  

x^9=1

+1=18

   =\left ( \cos(0)+i \sin(0) \right )^\frac{1}{9}

=\left ( \cos(2k\pi)+i \sin(2k\pi) \right )^{\frac{1}{9}},  k=0,1,2,...,8 8

  =\left ( \cos(\frac{2k\pi}{9})+i \sin(\frac{2k\pi}{9}) \right ),  k=0,1,2,...,8 8 [by De-Moivre's law]

  =\left ( \cos(\frac{2\pi}{9})+i \sin(\frac{2\pi}{9}) \right )^k ,  k=0,1,2,...,8 8 [by De-Moivre's law]

Let, \left ( \cos(\frac{2\pi}{9})+i \sin(\frac{2\pi}{9}) \right )=\alpha

\Rightarrow x=\alpha ^k,  k=0,1,2,...,8 8

Thus, the 9 roots of unity are given by 1,\alpha,\alpha^{2},\alpha^3,\cdots,\alpha^8 i.e.

1, \left ( \cos(\frac{2\pi}{9})+i \sin(\frac{2\pi}{9}) \right ) , \left ( \cos(\frac{4\pi}{9})+i \sin(\frac{4\pi}{9}) \right ) , \left ( \cos(\frac{6\pi}{9})+i \sin(\frac{6\pi}{9}) \right ) ,  \left ( \cos(\frac{8\pi}{9})+i \sin(\frac{8\pi}{9}) \right ), \left ( \cos(\frac{10\pi}{9})+i \sin(\frac{10\pi}{9}) \right ) ,  \left ( \cos(\frac{12\pi}{9})+i \sin(\frac{12\pi}{9}) \right ) ,  \left ( \cos(\frac{14\pi}{9})+i \sin(\frac{14\pi}{9}) \right ) , \left ( \cos(\frac{16\pi}{9})+i \sin(\frac{16\pi}{9}) \right )

b.) To find the 6th roots of unity:

We have,

x^6=-64

\Rightarrow x=\left ( -64 \right )^\frac{1}{6}

  =(-1\times 2^6)^\frac{1}{6}

  =(-1)^\frac{1}{6}\times \left ( 2^6 \right )^\frac{1}{6}

  =(-1)^\frac{1}{6}\times2

  =2(-1)^\frac{1}{6}

  =2(\cos(\pi)+i\sin(\pi))^\frac{1}{6}

  =2(\cos(\pi+2k\pi)+i\sin(\pi+2k\pi))^\frac{1}{6},\;k=0,1,2,3,4,5

  =2(\cos(\frac{\pi+2k\pi}{6})+i\sin(\frac{\pi+2k\pi}{6})),\;k=0,1,2,3,4,5   [by De-Moivre's law]

\Rightarrow x=2\cos(\frac{(2k+1)\pi}{6})+i\sin(\frac{(2k+1)\pi}{6})),\;k=0,1,2,3,4,5

Thus, the 6 roots of unity are given by 2(\cos(\frac{\pi}{6})+i\sin(\frac{\pi}{6})),\;2(\cos(\frac{3\pi}{6})+i\sin(\frac{3\pi}{6})),\;2(\cos(\frac{5\pi}{6})+i\sin(\frac{5\pi}{6})),\;2(\cos(\frac{7\pi}{6})+i\sin(\frac{7\pi}{6})),\;2(\cos(\frac{9\pi}{6})+i\sin(\frac{9\pi}{6})),\;2(\cos(\frac{11\pi}{6})+i\sin(\frac{11\pi}{6}))

i.e.  2(\frac{\sqrt{3}}{2}+i\frac{1}{2}),\;2(0+i),\;2(\frac{-\sqrt{3}}{2}+i\frac{1}{2}),\;2(\frac{-\sqrt{3}}{2}-i\frac{1}{2}),\;2(0-i),\;2(\frac{\sqrt{3}}{2}-i\frac{1}{2})

i.e. \left ( \sqrt{3}+i \right ),\;2i,\;\left ( -\sqrt{3}+i \right ),\;\left ( -\sqrt{3}-i \right ),\;-2i,\;\left ( \sqrt{3}-i \right )

NOTE:

According to the De-Moivre's formula all nth roots of a complex number r\left(\cos\left(\theta\right)+i\sin\left(\theta\right)\right) are given by

\sqrt[n]{r}\left(\cos\left(\frac{\theta+2\pi k}{n}\right)+i\sin\left(\frac{\theta+2\pi k}{n}\right)\right) , k=0,1,2,...,n-1